OpenStudy (anonymous):

The sum of two consecutive odd intergers is 236. What is the smallest interger? a. 115 b. 117 c. 119 d. 113

OpenStudy (anonymous):

x = 1st odd integer x + 2 = 2nd odd integer (x) + (x + 2) = 236 -- combine like terms 2x + 2 = 236 -- subtract 2 from both sides 2x = 236 - 2 2x = 234 -- divide by 2 on both sides x = 117 x + 2 = 117 + 2 = 119 your 2 integers are 117 and 119, with 117 being the smallest integer.

OpenStudy (anonymous):

Thanks kelliegirl33

OpenStudy (anonymous):

Thanks kelliegirl33

OpenStudy (anonymous):

Thanks kelliegirl33

OpenStudy (anonymous):

Thanks kelliegirl33